toms uncle owns a triangle piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?: toms uncle owns a triangle piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?
Answer by jojo14344(888)


Remember: P[T]=a+b+c ----------------------> EQN 1
where a=longest=highlight(x(yards))
b=2ndlongest=highlight(x-2(yards))
c=shortest=highlight((1/2)x+30(yards))
P=378 yards
Going back EQN 1,
378=x+x-2+(1/2)x+30
378=2(1/2)x+28
378-28=2(1/2)x
350=2(1/2)x
cross(350)140/cross(2(1/2))=cross(2(1/2))x/cross(2(1/2))
x=140yards LONGEST side (a)
140-2=138yards 2ND LONGEST (b)
(1/2)(140)+30=70+30=100yards SHORTEST (c)
Check via EQN 1:
378=140+138+100
378yards=378yards
Thank you,
Jojo